# Concentration of discrepancy–based approximate Bayesian computation via Rademacher complexity

### v4

@@ 2 Integral probability semimetrics (ips) and Rademacher complexity
(Summary–based distances). Classical summary–based abc employs distances among a finite set of summaries $f_{1},\ldots,f_{K}$ (e.g., Drovandi & Frazier, 2022). Interestingly, these distances can be recovered as a special case of mmd with suitably–defined kernel and, hence, are guaranteed to belong to the ips class. In particular, for a pre–selected vector of summaries $f(x)=[f_{1}(x),\ldots,f_{K}(x)]\in\mathcal{H}=\mathbb{R}^{K}$ equipped with the standard Euclidean norm $\|f(x)\|_{\mathcal{H}}^{2}=\left<f(x),f(x)\right>=f^{2}_{1}(x)+\cdots+f^{2}_{K}(x)$, one can define the kernel $k(x,x^{\prime})=\left<f(x),f(x^{\prime})\right>$ to obtain classical summary–based abc recasted within the mmd framework. As a result, popular kernels, such as the Gaussian one, relying on infinite feature maps can be interpreted as infinite dimensional versions of summary–based abc.
Although Examples 2.12.3 characterize the most popular instances of ips discrepancies employed in abc, it shall be emphasized that other interesting semimetrics belong to the ips class (e.g., Sriperumbudur et al., 2012; Birrell et al., 2022). Two relevant ones are the total variation (tv) and Kolmogorov–Smirnov distances discussed in the Supplementary Materials (see Examples A.1.1A.1.2).
While abc based on discrepancies, such as those presented above, overcomes the need to preselect summaries, it still requires the choice of a discrepancy. This motivates theory on the asymptotic properties of the induced abc posterior — under general discrepancies — for $n\rightarrow\infty$ and relevant thresholding schemes. More specifically, considering here a unified perspective which applies to the whole ips class, current theory focuses both on fixed $\varepsilon_{n}=\varepsilon$ settings, and also on $\varepsilon_{n}\rightarrow\varepsilon^{*}$ regimes, where $\varepsilon^{*}=\inf\nolimits_{\theta\in\Theta}\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ is the lowest attainable discrepancy between the assumed model and the data generating process (Jiang et al., 2018; Bernton et al., 2019; Nguyen et al., 2020; Frazier, 2020; Fujisawa et al., 2021). Under these settings, available theory investigates whether the $\varepsilon_{n}$–control on $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ established by rejection–abc translates into meaningful convergence results, and upper bounds on the rates of concentration for the sequence of discrepancy–based abc posteriors around the true data generating process $\mu^{*}$, both under correctly specified models where $\mu^{*}=\mu_{\theta^{*}}$ for some $\theta^{*}\in\Theta$, and in misspecified contexts where $\mu^{*}$ does not belong to $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$.
@@
$\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*}).$ (2.1)
To this end, available theoretical studies pre–assume suitable convergence results for the empirical measures along with specific concentration inequalities regulated by non–explicit sequences of control functions (see e.g., Assumptions 1–2 in Proposition 3 of Bernton et al. (2019), and Assumptions A1–A2 in Nguyen et al. (2020)). Alternatively, it is possible to directly require that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\rightarrow\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ almost surely as $n\rightarrow\infty$ (see Jiang et al., 2018; Nguyen et al., 2020, Theorems 1 and 2, respectively). However, as highlighted by the same authors, these conditions (i) can be difficult to verify for several discrepancies, (ii) do not allow to assess whether some of these discrepancies can achieve convergence and concentration uniformly over $\mathcal{P}(\mathcal{Y})$, and (iii) often yield bounds which hinder an in–depth understanding of the factors regulating these limiting properties. In addition, when the above assumptions are not met, the asymptotic behavior of the induced abc posteriors remains yet unexplored.
As a first step towards addressing the above issues, notice that the aforementioned results, when contextualized within the ips class, are inherently related to the richness of the known class of functions $\mathfrak{F}$ that identifies each $\mathcal{D}_{\mathfrak{F}}$; see Definition 2.1. Intuitively, if $\mathfrak{F}$ is too rich, it is always possible to find a function $f\in\mathfrak{F}$ yielding large discrepancies even when $\hat{\mu}_{z_{1:n}}$ and $\hat{\mu}_{y_{1:n}}$ are arbitrarily close to $\mu_{\theta}$ and $\mu^{*}$, respectively. Hence, $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ will remain large with positive probability, making the previous triangle inequality of limited interest, since a low $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ will not necessarily imply a small $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. In fact, in this context, it is clearly even not guaranteed that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ will converge to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. This suggests that the limiting properties of the abc posterior are inherently related to the richness of $\mathfrak{F}$. In Sections 3 and 6 we prove that this is the case when such a richness is measured via the notion of Rademacher complexity clarified in Definition 2.2; see also Chapter 4 in Wainwright (2019) for an introduction.
@@
Given a sample $x_{1:n}=(x_{1},\ldots,x_{n})$ of i.i.d. realizations from $\mu\in\mathcal{P}(\mathcal{Y})$, the Rademacher complexity $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ of a class $\mathfrak{F}$ of real–valued measurable functions is
@@
As is clear from the above definition, high values of the Rademacher complexity $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ mean that $\mathfrak{F}$ is rich enough to contain functions that can closely interpolate, on average, even full–noise vectors. Hence, if $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ is bounded away from zero for every $n$, then $\mathfrak{F}$ might be excessively rich for statistical purposes. Conversely, if $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ goes to zero as $n$ diverges, then $\mathfrak{F}$ is a more parsimonious class. Lemma 2.1 formalizes this intuition.
@@
Let $x_{1:n}$ be i.i.d. from some distribution $\mu\in\mathcal{P}(\mathcal{Y})$. Then, for any $b$–uniformly bounded class $\mathfrak{F}$, i.e., any class $\mathfrak{F}$ of functions $f$ such that $\|f\|_{\infty}\leq b$, any integer $n\geq 1$ and scalar $\delta\geq 0$, it holds that
$\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{\mu,n}(\mathfrak{F})+\delta\right]\geq 1-\exp(-n\delta^{2}/2b^{2}),$ (2.2)
and $\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\geq\mathfrak{R}_{\mu,n}(\mathfrak{F})/2-\sup_{f\in\mathfrak{F}}|\mathbb{E}(f)|/2n^{1/2}-\delta\right]\geq 1-\exp(-n\delta^{2}/2b^{2})$.
@@
Lemma 2.1 provides general bounds for the probability that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)$ takes values below a multiple and above a fraction of the Rademacher complexity of $\mathfrak{F}$. Recalling the previous discussion on the concentration of abc posteriors, this result is fundamental to study the convergence of both $\hat{\mu}_{z_{1:n}}$ and $\hat{\mu}_{y_{1:n}}$ to $\mu_{\theta}$ and $\mu^{*}$, respectively. Our theory in Section 3 proves that this is possible when $\mathfrak{R}_{n}(\mathfrak{F}):=\sup_{\mu\in\mathcal{P}(\mathcal{Y})}\mathfrak{R}_{\mu,n}(\mathfrak{F})$ goes to 00 as $n\rightarrow\infty$. According to Lemma 2.1, this is, in fact, a necessary and sufficient condition for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\rightarrow 0$ as $n\rightarrow\infty$, in $\mathbb{P}_{x_{1:n}}$–probability, uniformly over $\mathcal{P}(\mathcal{Y})$. See Section 6 for non–i.i.d. extensions.
3 Asymptotic properties of discrepancy–based abc posterior distributions
As anticipated in Section 2, the asymptotic results we derive in this article leverage Lemma 2.1 to connect the properties of the abc posterior under an ips discrepancy $\mathcal{D}_{\mathfrak{F}}$ with the behavior of the Rademacher complexity for the associated family $\mathfrak{F}$.
@@ 3 Asymptotic properties of discrepancy–based abc posterior distributions
Section 3.1 clarifies the importance of such a bridge by studying the limiting properties of the abc posterior when $\varepsilon_{n}=\varepsilon$ is fixed and $n\rightarrow\infty$, including in yet–unexplored situations where convergence guarantees for the empirical measures and for the discrepancy among such measures are not necessarily available. The theoretical results we derive in this setting suggest that the use of discrepancies whose Rademacher complexity decays to zero in the limit is a sufficient condition to obtain strong convergence guarantees. On the basis of these findings, Section 3.2 focuses on the regime $\varepsilon_{n}\rightarrow\varepsilon^{*}$ as $n\rightarrow\infty$, and derives unified and informative concentration bounds that crucially hold uniformly over $\mathcal{P}(\mathcal{Y})$, and are based on more constructive conditions than those employed in the available discrepancy–specific theory. More specifically, to prove our theoretical results in Sections 3.1 and 3.2, we will rely on some or all of the following assumptions:
@@
    there exist positive $L$ and $c_{\pi}$ such that, for $\bar{\varepsilon}$ small, $\pi(\{\theta{:}\mathcal{D}_{\mathfrak{F}}(\mu_{\theta}{,}\mu^{*})\leq\varepsilon^{*}{+}\bar{\varepsilon}\})\geq~{}c_{\pi}\bar{\varepsilon}^{L}$;
@@
    $\mathfrak{F}$ is $b$–uniformly bounded; i.e., there exists $b\in\mathds{R}^{+}$ such that $||f||_{\infty}{\leq b}$ for any $f\in\mathfrak{F}$;
@@
Condition (I) is the only assumption made on the data generating process and is present in, e.g., Nguyen et al. (2020) and in the supplementary materials of Bernton et al. (2019). Although the theory we derive in Section 6 relaxes (I) to study convergence and concentration also beyond the i.i.d. context, it shall be emphasized that some of the assumptions considered in the literature may not hold even in i.i.d. settings. Hence, an improved understanding of the abc properties under (I) is crucial to clarify the range of applicability and potential limitations of available existence theory under more complex, possibly non–i.i.d., regimes. In fact, as shown in Section 3.1, certain discrepancies may yield posteriors that are either not well–defined in the limit or lack strong convergence guarantees even in i.i.d. settings. Assumption (II) is standard in Bayesian asymptotics and abc theory (Bernton et al., 2019; Nguyen et al., 2020; Frazier, 2020), and requires that sufficient prior mass is placed on those parameter values yielding models $\mu_{\theta}$ close to $\mu^{*}$, under $\mathcal{D}_{\mathfrak{F}}$. Finally, (III)–(IV) provide conditions on $\mathfrak{F}$ to ensure that $\mathcal{D}_{\mathfrak{F}}$ yields meaningful concentration bounds, leveraging Lemma 2.1. Crucially, (III)–(IV) are made directly on the user–selected discrepancy $\mathcal{D}_{\mathfrak{F}}$, rather than on the data generating process or on the statistical model, thereby providing general constructive conditions that are more realistic to verify in practice compared to regularity assumptions on the unknown $\mu^{*}$ or on $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, while guiding the choice of $\mathcal{D}_{\mathfrak{F}}$.
@@
As clarified in Section 3.3, conditions (III)–(IV) can be broadly verified leveraging known bounds on the Rademacher complexity, and hold for the relevant ips discrepancies in Examples 2.12.3 under either no additional constraints or for suitable spaces $\mathcal{Y}$ (e.g., bounded ones); see Examples 3.13.3 and refer to the Supplementary Materials for the analysis of other relevant ips discrepancies. Moreover, these conditions are generally not stronger than those implicit in current theory on discrepancy–based abc, and are inherently related with the notion of uniform Glivenko–Cantelli classes (see e.g., Wainwright, 2019, Chapter 4), arguably a minimal sufficient requirement to establish convergence and concentration properties that, unlike those currently derived, hold uniformly over $\mathcal{P}(\mathcal{Y})$. It shall be also emphasized that the convergence theory derived in Section 3.1 does not necessarily require a vanishing Rademacher complexity to obtain informative results. However, it also clarifies that, when this condition is verified, the limiting pseudo–posterior inherits the same threshold–control on $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ as the one established by the abc procedure on $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$.
@@
As anticipated above, Sections 3.13.2 consider the two scenarios where the sample size $n$ diverges to infinity, while the abc threshold $\varepsilon_{n}$ is either fixed to $\varepsilon$ or progressively shrinks to $\varepsilon^{*}=\inf\nolimits_{\theta\in\Theta}\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. If the model is well–specified, i.e., $\mu^{*}=\mu_{\theta^{*}}$ for some, possibly not unique, $\theta^{*}\in\Theta$, then $\varepsilon^{*}=0$. Although the regime characterized by $\varepsilon_{n}\to 0$ with fixed $n$ is also of interest, this setting can be readily addressed under the unified ips class via a direct adaptation of available results in, e.g., Jiang et al. (2018) and Bernton et al. (2019); see also Miller and Dunson (2019) for results in the context of coarsened posteriors under the regime with $n\to\infty$ and fixed $\varepsilon_{n}=\varepsilon$, related to those in Jiang et al. (2018).

### v5

@@ 2 Integral probability semimetrics and Rademacher complexity
(Summary-based distances). Classical summary-based abc employs distances among a finite set of summaries $f_{1},\ldots,f_{K}$ (e.g., Drovandi and Frazier, 2022). Interestingly, these distances can be recovered as a special case of mmd with suitably-defined kernel and, hence, are guaranteed to belong to the ips class. In particular, for a pre-selected vector of summaries $f(x)=[f_{1}(x),\ldots,f_{K}(x)]\in\mathcal{H}=\mathbb{R}^{K}$ equipped with the standard Euclidean norm $\|f(x)\|_{\mathcal{H}}^{2}=\left<f(x),f(x)\right>=f^{2}_{1}(x)+\cdots+f^{2}_{K}(x)$, one can define the kernel $k(x,x^{\prime})=\left<f(x),f(x^{\prime})\right>$ to obtain classical summary-based abc recasted within the mmd framework. As a result, popular kernels, such as the Gaussian one, relying on an infinite dimensional feature space can be interpreted as a limiting version of summary-based abc.
Although Examples 2.22.4 characterize the most popular instances of ips discrepancies employed in abc, it shall be emphasized that other interesting semimetrics belong to the ips class (e.g., Sriperumbudur et al., 2012; Birrell et al., 2022). Two relevant ones are the total variation (tv) and Kolmogorov–Smirnov distances discussed in the Supplementary Material.
While abc based on discrepancies, such as those presented above, overcomes the need to pre-select summaries, it still requires the choice of a discrepancy. This motivates theory on the asymptotic properties of the induced abc posterior, under general discrepancies, for $n\rightarrow\infty$ and relevant thresholding schemes. Considering here a unified perspective which applies to the whole ips class, the current theory focuses both on fixed $\varepsilon_{n}=\varepsilon$ settings, and also on $\varepsilon_{n}\rightarrow\varepsilon^{*}$ regimes, where $\varepsilon^{*}=\inf\nolimits_{\theta\in\Theta}\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ is the lowest attainable discrepancy between the assumed model and the data generating process (Jiang, Wu and Wong, 2018; Bernton et al., 2019; Nguyen et al., 2020; Frazier, 2020; Frazier, Robert and Rousseau, 2020; Fujisawa et al., 2021). Under these settings, available theory investigates whether the $\varepsilon_{n}$-control on $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ established by rejection-abc translates into meaningful convergence results, and upper bounds on the rates of concentration for the sequence of discrepancy-based abc posteriors around the true data generating process $\mu^{*}$, both under correctly specified models where $\mu^{*}=\mu_{\theta^{*}}$ for some $\theta^{*}\in\Theta$, and in misspecified contexts where $\mu^{*}$ does not belong to $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$.
@@
$\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*}).$ (1)
To this end, available theoretical studies pre-assume suitable convergence results for the empirical measures along with specific concentration inequalities regulated by non-explicit sequences of control functions; see Assumptions 1–2 in Proposition 3 of Bernton et al. (2019), and Assumptions A1–A2 in Nguyen et al. (2020). Alternatively, it is possible to directly require that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\rightarrow\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ almost surely as $n\rightarrow\infty$ (see Jiang, Wu and Wong, 2018; Nguyen et al., 2020, Theorems 1 and 2, respectively). However, as highlighted by the same authors, these conditions (i) can be difficult to verify for several discrepancies, (ii) do not allow to assess whether some of these discrepancies can achieve convergence and concentration uniformly over $\mathcal{P}(\mathcal{Y})$, and (iii) often yield bounds which hinder an in-depth understanding of the factors regulating these limiting properties. In addition, when the above assumptions are not met, the asymptotic behavior of abc posteriors remains yet unexplored.
As a first step towards addressing the above issues, notice that the aforementioned results, when contextualized within the ips class, are inherently related to the richness of the known class of functions $\mathfrak{F}$ that identifies each $\mathcal{D}_{\mathfrak{F}}$; see Definition 2.1. Intuitively, if $\mathfrak{F}$ is too rich, it is always possible to find a function $f\in\mathfrak{F}$ yielding large discrepancies between the empiricals and the corresponding truths even when $\hat{\mu}_{z_{1:n}}$ and $\hat{\mu}_{y_{1:n}}$ are arbitrarily close to $\mu_{\theta}$ and $\mu^{*}$, respectively. Hence, $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ will remain large with positive probability, making the triangle inequality in (1) of limited interest, since a low $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ will not necessarily imply a small $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. In fact, in this context, it is clearly even not guaranteed that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ will converge to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. This suggests that the limiting properties of the abc posterior are inherently related to the richness of $\mathfrak{F}$. In Section 3 we prove that this is the case when such a richness is measured through the notion of Rademacher complexity clarified in Definition 2.5; see also Chapter 4 in Wainwright (2019) for an introduction.
@@
Given a vector $x_{1:n}=(x_{1},\ldots,x_{n})$ of i.i.d. random variables with distribution $\mu\in\mathcal{P}(\mathcal{Y})$, the Rademacher complexity $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ of a class $\mathfrak{F}$ of real-valued measurable functions is
@@
As is clear from the above definition, high values of the Rademacher complexity $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ mean that $\mathfrak{F}$ is rich enough to contain functions that can closely capture, on average, even the behavior of full-noise vectors. Hence, if $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ is bounded away from zero for every $n$, then $\mathfrak{F}$ might be excessively rich for statistical purposes. Conversely, if $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ goes to zero as $n$ diverges, then $\mathfrak{F}$ is a more parsimonious class. Lemma 2.6 formalizes this intuition.
@@
Let $x_{1:n}$ be i.i.d. from some distribution $\mu\in\mathcal{P}(\mathcal{Y})$. Then, for any $b$-uniformly bounded class $\mathfrak{F}$, i.e., any class $\mathfrak{F}$ of functions $f$ such that $\|f\|_{\infty}\leq b$, any integer $n\geq 1$ and scalar $\delta\geq 0$, it holds that
$\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{\mu,n}(\mathfrak{F})+\delta\right]\geq 1-\exp(-n\delta^{2}/2b^{2}),$ (2)
and $\mathbb{P}_{x_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\geq\mathfrak{R}_{\mu,n}(\mathfrak{F})/2-\sup_{f\in\mathfrak{F}}|\mathbb{E}(f)|/2n^{1/2}-\delta]\geq 1-\exp(-n\delta^{2}/2b^{2})$.
@@
Lemma 2.6 provides bounds for the probability that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)$ takes values below a multiple and above a fraction of the Rademacher complexity of $\mathfrak{F}$. Recalling the previous discussion on concentration of abc posteriors, this result is crucial to study the convergence of both $\hat{\mu}_{z_{1:n}}$ and $\hat{\mu}_{y_{1:n}}$ to $\mu_{\theta}$ and $\mu^{*}$, respectively. Our theory in Section 3 proves that this is possible when $\mathfrak{R}_{n}(\mathfrak{F}):=\sup_{\mu\in\mathcal{P}(\mathcal{Y})}\mathfrak{R}_{\mu,n}(\mathfrak{F})$ goes to 00 as $n\rightarrow\infty$. According to Lemma 2.6, this is a necessary and sufficient condition for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\rightarrow 0$ as $n\rightarrow\infty$, in $\mathbb{P}_{x_{1:n}}$–probability, uniformly over $\mathcal{P}(\mathcal{Y})$. See Appendix C in the Supplementary Material for extensions to non-i.i.d. settings.
3 Asymptotic properties of discrepancy-based abc posterior distributions
@@
As anticipated in Section 2, the asymptotic results we derive leverage Lemma 2.6 to connect the properties of the abc posterior under an ips discrepancy $\mathcal{D}_{\mathfrak{F}}$ with the behavior of the Rademacher complexity for the associated family $\mathfrak{F}$. Section 3.1 clarifies the importance of this bridge by studying the limiting properties of the abc posterior when $\varepsilon_{n}=\varepsilon$ is fixed and $n\rightarrow\infty$, including in yet-unexplored situations where convergence guarantees for the empirical measures and for the discrepancy among such measures are not necessarily available. The theoretical results we derive in this case suggest that the use of discrepancies whose Rademacher complexity decays to zero in the limit is a sufficient condition to obtain strong convergence guarantees. On the basis of these findings, Section 3.2 focuses on the regime $\varepsilon_{n}\rightarrow\varepsilon^{*}=\inf\nolimits_{\theta\in\Theta}\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ as $n\rightarrow\infty$, and derives unified and informative concentration bounds that hold uniformly over $\mathcal{P}(\mathcal{Y})$ and are based on more constructive conditions than those employed in the available discrepancy-specific theory. More specifically, to prove our theoretical results in Sections 3.13.2, we rely on some or all of the following assumptions:
@@
    there exist $L,c_{\pi}\in\mathds{R}^{+}$ such that, for $\bar{\varepsilon}$ small, $\pi(\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}\})\geq c_{\pi}\bar{\varepsilon}^{L}$;
@@
    $\mathfrak{F}$ is $b$-uniformly bounded; i.e., there exists $b\in\mathds{R}^{+}$ such that $||f||_{\infty}{\leq b}$ for any $f\in\mathfrak{F}$;
@@
Condition (I) is the only assumption made on the data generating process and is present in, e.g., Nguyen et al. (2020) and in the supplementary materials of Bernton et al. (2019). Although the theory we derive in Appendix C of the Supplementary Material relaxes (I) to study convergence and concentration also beyond the i.i.d. context, it shall be emphasized that some of the assumptions considered in the literature may not hold even in i.i.d. settings. Hence, an improved understanding of the abc properties under (I) is crucial to clarify the range of applicability and potential limitations of available existence theory under more complex, possibly non-i.i.d. regimes. In fact, as shown within Section 3.1, certain discrepancies may yield posteriors that are either not well-defined in the limit or lack strong convergence guarantees even in i.i.d. settings. Assumption (II) is standard in Bayesian asymptotics and abc theory (Bernton et al., 2019; Nguyen et al., 2020; Frazier, 2020), and requires that sufficient prior mass is placed on those parameter values yielding models $\mu_{\theta}$ close to $\mu^{*}$, under $\mathcal{D}_{\mathfrak{F}}$. Finally, (III)–(IV) provide conditions on $\mathfrak{F}$ to ensure that $\mathcal{D}_{\mathfrak{F}}$ yields meaningful concentration bounds, leveraging Lemma 2.6. Crucially, (III)–(IV) are made directly on the user-selected discrepancy $\mathcal{D}_{\mathfrak{F}}$, rather than on the data generating process or on the statistical model, thereby providing general constructive conditions that are more realistic to verify in practice compared to regularity assumptions on the unknown $\mu^{*}$ or on $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, while guiding the choice of $\mathcal{D}_{\mathfrak{F}}$.
@@
Notice that (III)–(IV) depend, in principle, on $\mathcal{P}(\mathcal{Y})$. Nonetheless, as clarified within Section 3.3, available bounds on the Rademacher complexity ensure that these two conditions hold for the relevant ips discrepancies in Examples 2.22.4 under either no constraints on $\mathcal{P}(\mathcal{Y})$ — thereby circumventing such a dependence — or for suitable spaces $\mathcal{Y}$ (e.g., bounded ones). Unlike for constraints on the unknown $\mu^{*}$ and on $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, conditions on $\mathcal{Y}$ can be verified in practice; see Examples 3.53.7, and refer to the Supplementary Material for the analysis of other relevant ips discrepancies. Moreover, Assumptions (III) and (IV) are generally not stronger than those implicit in the current theory on discrepancy-based abc, and are inherently related with the notion of uniform Glivenko–Cantelli classes (see e.g., Wainwright, 2019, Chapter 4), arguably a minimal sufficient requirement for establishing convergence and concentration properties that, unlike those currently derived, hold uniformly over $\mathcal{P}(\mathcal{Y})$. This latter point is further clarified in Propositions 4.3 and 4.4, which state concentration results for mmd with unbounded kernel in $\mathbb{R}^{d}$ (which also includes abc with unbounded summary statistics) and Wasserstein-1 distance in $\mathbb{R}^{d}$, respectively. Although these two discrepancies meet (III)–(IV) for suitable constraints on $\mathcal{Y}$, in general unbounded $\mathcal{Y}$ settings these conditions are no more guaranteed for such discrepancies. In this case, Propositions 4.3 and 4.4 clarify that concentration properties can still be derived, but at the expense of additional constraints on $\mu^{*}$ and $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, which yield results that hold uniformly only in a suitable subset of $\mathcal{P}(\mathcal{Y})$, rather than over the entire $\mathcal{P}(\mathcal{Y})$. It shall be also emphasized that the convergence theory derived in Section 3.1 does not necessarily require a vanishing Rademacher complexity to obtain informative results. However, it also clarifies that, when this condition is verified, the limiting pseudo-posterior inherits the same threshold-control on $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ as the one established by rejection abc on $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$.
@@
As anticipated above, Sections 3.13.2 consider the two scenarios where the sample size $n$ diverges to infinity, while the abc threshold $\varepsilon_{n}$ is either fixed at $\varepsilon$ or progressively shrinks to $\varepsilon^{*}=\inf\nolimits_{\theta\in\Theta}\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. If the model is well-specified, i.e., $\mu^{*}=\mu_{\theta^{*}}$ for some, possibly not unique, $\theta^{*}\in\Theta$, then $\varepsilon^{*}=0$. Although the regime characterized by $\varepsilon_{n}\to 0$ with fixed $n$ is also of interest, this setting can be readily addressed under the unified ips class via a direct adaptation of available results in, e.g., Jiang, Wu and Wong (2018) and Bernton et al. (2019); see also Miller and Dunson (2019) for results in the context of coarsened posteriors under the regime with $n\to\infty$ and fixed $\varepsilon_{n}=\varepsilon$, related to those in Jiang, Wu and Wong (2018).